Hello! I've tried this website once and it really worked! So here's my question involving radical expressions:
Instruction(s): Find all real solutions in each equation.
1.) √(x-3)+1=√(x+2)
So I have tried to go to other websites such as Mathway, which had an answer of 7, but unfortunately, you have to pay to have the solutions.(can you please add solutions) Please answer back because I really need this tomorrow for school. Thanks a lot! More Power to you and algebra.com!

Answer by ewatrrr(24785)

 
Hi, Work-Up Previously Posted
√(x-3)+1=√(x+2) || Squaring boths sides of the EQ
(√(x-3)+1)(√(x-3)+1) = x+2
x-3 + 2√(x-3)+ 1 = x+2
2√(x-3) = 4
√(x-3) = 2 || Squaring boths sides of the EQ
x-3 = 4
x = 7